# Single attosecond XUV pulse source via light-wave controlled relativistic laser-plasma interaction: Thomson Back Scattering Scheme

### v1

@@ Abstract
Reflecting light off a mirror moving near light speed, as first envisioned by Einstein, offers a powerful method for generating bright, ultrashort pulses in the extreme ultraviolet range. Recent breakthroughs show that dense relativistic electron mirrors can be created by striking a nanometre-scale foil with a high-intensity, sharp-front laser pulse, forming a single relativistic electron sheet (RES). This RES coherently reflects and upshifts a counter-propagating laser beam from the infrared to the extreme ultraviolet with efficiency exceeding incoherent scattering by over several orders of magnitude. Here we demonstrate that optimizing the drive laser waveform can reliably produce a single RES, leading to generation of isolated attosecond pulses enhancing both intensity and temporal compression of the back reflected light in a controlled manner. Simulations reveal that tuning parameters like timing delay enables control over the amplitude, duration, and bandwidth of the resulting attosecond Thomson backscattering (TBS) pulse. Together, these advances meet key experimental challenges and pave the way for compact, tunable sources of isolated attosecond pulses for probing ultrafast phenomena.
@@ I Introduction
To generate isolated as pulses via CTS, the interaction must be confined to a single reflection event from a relativistically moving electron bunch [25]. This requires the formation of a solitary, ultra-thin, and coherent electron mirror propagating at relativistic velocities, with which a counter-propagating laser pulse can interact. A well-established method for producing high-quality relativistic electron bunches is laser wakefield acceleration (LWFA), where intense, ultra-short laser pulses interact with underdense plasma. In this regime, electrons are accelerated to relativistic energies and can subsequently interact with a counter-propagating laser pulse [26, 27, 28, 29]. While LWFA can produce highly energetic electron bunches, achieving the high peak electron densities required for forming a relativistic electron sheet (RES) remains a significant challenge. One promising approach involves the interaction of intense laser pulses with nanometer-scale solid-density foil targets. This method can generate relativistic electron bunches with optimized density and energy distributions, enabling the production of intense attosecond pulses. In such a backscattering scheme, two counter-propagating laser pulses—a drive and a source—interact with an ultra-thin solid foil [23, 24]. The drive pulse must be sufficiently intense to expel electrons from the foil and accelerate them to relativistic velocities before the target disintegrates. The resulting overdense, nanometer-scale RES, when illuminated by the counter-propagating few-cycle source pulse, emits spatially coherent, Doppler-upshifted radiation with sub-femtosecond temporal duration. Although, the use of multiple RESs for high harmonic generation (HHG) has been demonstrated experimentally with multi-cycle, linearly polarized laser pulses [30], the generation of a single dense RES capable of producing isolated attosecond pulses remains an open challenge.
@@ II Influence of drive laser waveform on TBS condition in laser-foil interaction
To enable the formation of an optimal Relativistic Electron Sheet (RES), the incident drive laser must exert a sufficiently strong ponderomotive force to expel plasma electrons from the ion background and accelerate them coherently in the forward direction. This requires the laser to generate a longitudinal push force capable of compressing electrons into a sub-skin-depth layer, ultimately overcoming the electrostatic restoring force of the ions. The necessary condition for this ”snowploughing” [12, 31] effect is that the normalized amplitude of the drive laser satisfies, $a_{0}^{d}>\omega_{p}^{2}d/2c\omega_{L}$, where $\omega_{p}=\sqrt{n_{e}e^{2}/\epsilon_{0}m_{e}}$ is the plasma frequency, $d$ is the initial thickness of the plasma foil, $\omega_{L}$ is the drive laser frequency and $c$ is the speed of light [32].
@@
Figure 1: Influence of drive pulse temporal envelope and its waveform on the formation of the electron mirror and backscattered signal. The $40$nm thick plasma slab (located from $6$ to $6.05\lambda_{L}$) is irradiated by the drive pulse possessing different waveform and temporal envelope (a) top-hat, (b) Gaussian, and (c) optimal dual-color Gaussian (with $P=0.4$) pulse, which causes various temporal evolution of electron number density. In all cases the drive beam with $a_{0}^{d}=20$ propagates from left to right and is incident normally on the $40$nm thick pre-ionized plasma foil at $6T_{L}$. (d) compares the temporal waveforms of the single ($1\omega$) and two-color ($1\omega+2\omega$) drive Gaussian pulses with end-to-end $12T_{L}$ temporal duration and polarization along y-direction. The inset at the lower left-hand side enlarges the waveforms from $4$ to $5T_{L}$, highlighting the steepness of accelerating two-color field. The counter propagating source pulse with duration $6T_{L}$ and peak amplitude of $a_{0}=0.1$ interacts with the electron sheet and reflects partially. The squared field of the normalized TBS signals are shown in (e), indicating the generation of a single $110$as pulse (in FWHM) in case of $(1\omega+2\omega)$ drive beam. The duration of the TBS signal, result from interaction with the produced RES in panel (a) is $380$as. This pulse is scaled by a factor of $10$ (in blue). The reflected pulses are temporally shifted for better comparison. The corresponding frequency spectra of the TBS radiations are illustrated in panel (f). The AP generated by interaction of dual-color drive beam and foil target is composed of a narrow bandwidth spectrum peaked at $26\omega_{L}$.
@@
To gain a comprehensive understanding of how the temporal waveform of the drive laser pulse influences the interaction dynamics with a dense plasma foil, we carry out 2D plane-wave particle-in-cell (PIC) simulations [33]. We examine the spatio-temporal evolution of the electron number density $n_{e}(x,t)$ (normalized by critical density $n_{c}$) as shown in Figure  1(a). This figure illustrates the interaction of a sharp-front, top-hat laser pulse—modeled as a sinusoidal wave—with a nanometric solid-density foil. The incident laser is linearly polarized (electric field along the y-axis) and propagates along the x-axis (left to right) with normalized electric field amplitude of $a_{0}^{d}=eE_{0}^{d}/mc\omega_{L}=20$ corresponding to a peak intensity of $8.6\times 10^{20}$/cm2 for a central wave length $\lambda_{L}=800$nm. The laser pulse impinges normally on a $40$nm -thick, sharp-edged fully ionized Carbon plasma foil initially located at $x=6\lambda_{L}$, with an initial electron density of $n_{e}=21n_{c}$. This configuration ensures a relativistic near-critical density regime, characterized by the similarity parameter $S=n_{e}/a_{d}n_{c}\approx 1$ [34], which is optimal for efficient electron blowout and RES formation. Here, the critical density is defined as $n_{c}=mc\omega^{2}_{L}/e^{2}=1.7\times 10^{21}$cm-3 where $m$ denote the electron mass. The initial temperature of the plasma electrons is set as $T_{e}=100$eV. The size of each spatial grid of the simulation box is $\Delta x=\lambda_{L}/4000$, in which there are $500$ macro-particles of each species.
@@
As the sharp-fronted drive pulse reaches the target at $t=6T_{L}$, the radiation pressure at the leading edge of the pulse rapidly exceeds the electrostatic binding forces between electrons and ions. This results in a prompt and coherent expulsion of electrons in the forward direction, parallel to the laser propagation axis. The expelled electrons are compressed into a dense, ultra-thin relativistic electron sheet (RES) that is accelerated synchronously as is evident in Figure 1(a). Although such an approach using non-adiabatic sharp rising pulse has been investigated in numerous occasions for generating RES accessing such pulses are extremely complex and challenging. Considering the key role of pulse front steepness, plasma-based pulse front steepening was proposed in [35] and implemented in ion acceleration experiments [36]. In such experiments a separate low density target is designed prior to the main target, where relativistic non-linearities introduced by this near critical density plasma modifies the front edge of the pulse. This approach ultimately leads to a laser pulse, which is spatio-temporally compressed [37]. Once such pulses are generated single RES can be formed [38] and then in combination with double layer targets [39, 40] there are different propositions to boost the longitudinal Lorentz factor $\gamma_{x}$ and enhance the backreflected as pulse. Nontheless, such methods based on nonlinear pulse shaping in plasma medium involves higher order of experimental complexity.
@@
A typical currently available, ultrashort high power laser system emanates pulses with Gaussian temporal envelop on target. Figure 1(b) presents the electron density evolution when the same plasma foil is irradiated by a Gaussian-shaped laser pulse with identical peak field amplitude $a_{0}^{d}=20$. Despite having the same peak intensity, the gradual rise of the Gaussian envelope fails to deliver a sufficiently strong ponderomotive force at the pulse front. As a result, the electrons are not coherently swept into a single sheet but instead experience multiple oscillations within the laser field. This leads to incomplete electron blowout, in contrast to the case presented in Figure 1(a), and the plasma surface becomes distorted and modulated, rather than forming a clean RES. The electrons remain partially bound and oscillate around the ion background, resulting in a disordered plasma structure with reduced coherence and acceleration efficiency. It is also evident in Figure 1(b), that there are signatures of partial electron ejection on the back side of the irradiated target, leading to multiple potential electron bunches of varying strength. This comparison highlights the critical role of the pulse temporal profile—particularly the sharpness of the leading edge—in determining the efficiency of RES formation. A steep front ensures a rapid onset of radiation pressure, enabling synchronized electron expulsion, while a smooth Gaussian front leads to phase mixing and multiple bunch emission leading to their structural degradation.
In Figure 1(a), as the drive pulse with sharp front evolves to the target, rapidly rising radiation pressure at the pulse front quickly surpasses the electrostatic binding forces between plasma electrons and ions. This results in the coherent expulsion of electrons in the forward direction (along the laser propagation axis), leading to the formation of a dense, ultra-thin RES. The sharp temporal edge of the pulse ensures that the ponderomotive force acts collectively and synchronously on the entire electron population, enabling efficient RES formation.
In contrast, as shown in Figure 1(b), when a Gaussian-shaped drive pulse with the same peak field amplitude $a_{0}^{d}=20$ is used, the gradual rise of the pulse envelope fails to deliver a sufficiently strong push force at the leading edge. As a result, the plasma electrons are not expelled as a single coherent bunch but instead undergo multiple oscillations within the laser field. This leads to partial electron displacement, surface deformation, and the formation of a distorted plasma structure, where electrons oscillate around the ion background rather than forming a clean RES.
@@
This distinction in interaction dynamics is particularly relevant in the context of coherent Thomson backscattering for generation of as pulses. In the experiment reported by Kiefer et. al. [30], a $10$nm diamond-like carbon (DLC) was irradiated by a Gaussian-shaped drive pulse with intensity exceeding $1\times 10^{20}$cm-2. Due to the smooth rising edge of the Gaussian pulse, the initial few laser cycles were insufficiently intense to fully overcome the binding forces within the foil. Consequently, only a fraction of the electrons were expelled during each half-cycle of the laser field, resulting in the formation of multiple relativistic electron layers, or relativistic electron mirrors (REMs), at the rear side of the target. These REMs, being only partially reflective, interact with a counter-propagating probe pulse of intensity $\sim 1\times 10^{15}$cm-2. As the probe reflects off successive REMs, it accumulates multiple Doppler upshifts, producing a reflected spectrum rich in high-order harmonics of the probe’s fundamental frequency. This mechanism stands in contrast to the interaction driven by a sharp-edged pulse, which is capable of expelling all plasma electrons in a single cycle, thereby forming a single, highly reflective REM. Such a configuration is more favorable for generating isolated attosecond bursts and achieving spectrally cleaner harmonic emission.
@@
Since the RES formation is directly governed by the drive pulse field, there exists a strong correlation between the waveform of the drive pulse and the coherent electron dynamics responsible for RES generation [23, 24, 41]. Here, we demonstrate that the laser field waveform can be shaped optimally by the addition of a second harmonic component to a Gaussian drive pulse to excite plasma electron dynamics, that enables the formation and acceleration of a single, dense relativistic electron sheet. The significance of drive pulse waveform engineering has been widely explored in various high-field plasma interaction regimes, including laser wakefield acceleration [42] and attosecond XUV pulse generation via high harmonic emission from plasma surfaces [43, 44, 7, 45, 46, 47]. These studies consistently highlight the importance of sub-cycle control over the laser field to manipulate electron trajectories and optimize emission characteristics.
@@
As indicated by the arrow in Figure 1(c), this two-color linearly polarized drive configuration effectively excites the plasma electrons into a coherent structure. Figure 1(d) illustrates the two drive pulses employed to initiate interaction different drive pulse envelopes: top-hat (blue), Gaussian (in black marked $1\omega$), and two-color Gaussian (in magenta marked $1\omega+2\omega$). Importantly, in this set of simulations, the total energy of the drive pulse incident on the target is kept constant across all three cases—top-hat, Gaussian, and two-color Gaussian. However, the temporal waveform of each pulse leads to distinct energy distributions at the target surface, which in turn results in markedly different electron dynamics and RES formation behavior.
Once a single forward propagating RES is formed, a cross-polarized, counter-propagating source pulse is introduced at a carefully timed delay relative to the drive pulse. This source pulse interacts with the RES at normal incidence and is reflected, undergoing a Doppler frequency upshift and temporal compression into a single attosecond pulse (AP). The efficacy of RES for providing optimal conditions for Thomson backscattering (TBS) of a linearly polarized Gaussian source pulse (electric field along the z-axis), with normalized amplitude with $a_{0}=0.1$ and a temporal duration of $4$fs (corresponding to $1.5T_{L}$ in FWHM of intensity) is summarized in Figure 1(e-f). This correlation between the quality of RES and back reflected pulse, is quantitatively illustrated in Figure 1(e), which compares the temporal profiles of TBS pulses generated by RESs formed under different drive pulse temporal waveforms. The two-color drive pulse produces a compressed AP with a duration of (intensity FWHM) $110$as and a peak intensity of $a_{0}^{2}\approx 1.2\times 10^{-3}$, significantly outperforming the weaker and longer AP ($380$as) generated by even the ideal top-hat pulse. This clearly demonstrates the enhanced efficiency and coherence enabled by waveform shaping. The temporal profile for the Gaussian case does not lead to single AP formation.
The corresponding frequency spectra presented in Figure 1(f) further reinforce this conclusion. The ideally top-hat drive pulse is plotted in blue curve and bicolor Gaussian pulse in magenta curve. The RES formed by the two-color drive pulse reaches a higher velocity, resulting in a larger central frequency shift of the reflected spectrum up to $26\omega_{L}$ and a broader bandwidth compared to the sharp rising pulse from case, ideal for shorter AP. In contrast, in case of Gaussian drive (black curve in Figure 1(f)), most of the scattered energy is concentrated in low order harmonics induced in the reflected pulse, in line with the observation discussed in Figure 1(b).
These results underscore the critical role of multi-frequency waveform synthesis in controlling RES formation and optimizing attosecond pulse generation. A deeper analysis of this mechanism is presented in Section III, where the dynamics of single REM formation and its impact on coherent backscattering are discussed in detail.
@@
Figure 2: Optimal generation of single RES with waveform-controlled laser driver. (a) The longitudinal velocity of the ejected electron beam, averaged over the bunch $<\!\beta_{x}\!>$ as a function of $P$ and $\Delta\phi$. (b) The waveform dependent maximum energy of accelerated electrons in the bunch represented by the acquired cutoff energy $\varepsilon_{c}$ in ($P,\Delta\phi$)-space. The dashed encircled region in (a) and (b) denotes the parameter space, in which a single RES is produced and accelerated. The density profile of the electron beam after traveling $5\lambda_{L}$ to the location of detector $(x=11\lambda_{L})$ are plotted for two optimal waveforms of the drive laser (c) $P=0.4$ with $\Delta\phi=0$, and (d) $P=0.6$ with $\Delta\phi=-\pi/3$, in which a single electron beam with attosecond duration is emitted. (e) shows the spatio-temporal evolution of thin foil electron density (red to yellow colormap) and drive pulse field (green to violet colormap) for a single color Gaussian driver. (f) The spacetime evolution of the laser field and electron density for single RES regime as well as the temporal evolution of phasespace parameters of electron bunches. The time axis implies the simulation time (see Fig. 1(b) and (c)) during interaction of drive beam, synthesized by (e) only fundamental frequency ($1\omega$) or (f) by combination of in phase frequencies $(1\omega+2\omega)$ with energy ratio $P=0.4$. The dashed lines in (e) and (f) represent trajectory of the representative electrons which are subject to the normalized drive field $B_{z}$. (g) and (h) show temporal variation of the longitudinal and transverse components of the velocity of the representative electrons (black curve for $1\omega$ and magenta for $1\omega+2\omega$ field). (i) Temporal evolution of bunch peak density and (j) thickness during interaction interval with counter streaming source pulse with the RES produced two-color Gaussian driver beam with $P=0.4$ with $\Delta\phi=0$. The FWHM of the bunch is calculated by fitting a Gaussian function on the electron density profile. Greater longitudinal velocity of RES during interaction phase accounts for larger compression of the AP in case of synthesized pulse.

### v2

@@ Abstract
Reflecting light off a mirror moving near light speed offers a powerful method for generating bright, ultrashort pulses in the extreme ultraviolet range. Several investigations show that dense relativistic electron mirrors can be created by striking a nanometre-scale foil with a high-intensity, sharp-front laser pulse, forming a single relativistic electron sheet (RES). This RES coherently reflects and upshifts a counter-propagating laser beam from the infrared to the extreme ultraviolet with efficiency exceeding incoherent scattering by over several orders of magnitude. Here we demonstrate that optimizing the drive laser waveform can reliably produce a single RES, leading to generation of isolated attosecond pulses enhancing both intensity and temporal compression of the back reflected light in a controlled manner. Simulations reveal that tuning parameters like timing delay enables control over the amplitude, duration, and bandwidth of the resulting attosecond Thomson backscattering pulse. Together, these advances meet key experimental challenges and pave the way for compact, tunable sources of isolated attosecond pulses for probing ultrafast phenomena.
@@ I Introduction
To generate isolated as pulses via CTS, the interaction must be confined to a single reflection event from a relativistically moving electron bunch [25]. This requires the formation of a solitary, ultra-thin, and coherent electron mirror propagating at relativistic velocities, with which a counter-propagating laser pulse can interact. A well-established method for producing high-quality relativistic electron bunches is laser wakefield acceleration (LWFA), where intense, ultra-short laser pulses interact with underdense plasma. In this regime, electrons are accelerated to relativistic energies and can subsequently interact with a counter-propagating laser pulse [26, 27, 28, 29]. Although LWFA can produce highly energetic electron bunches, achieving the high peak electron densities required for the formation of a RES remains a significant challenge. One promising approach, not based on LWFA, involves the interaction of intense laser pulses with nanometer-scale solid-density foil targets. This method can generate relativistic electron bunches with optimized density and energy distributions, enabling the production of intense attosecond pulses. In such a backscattering scheme, two counter-propagating laser pulses interact with an ultra-thin solid foil [23, 24]. The drive pulse must be sufficiently intense to expel electrons from the foil and accelerate them to relativistic velocities before the target disintegrates. The resulting overdense, nanometer-scale RES, when illuminated by the counter-propagating few-cycle source pulse, emits spatially coherent, Doppler-upshifted radiation with sub-femtosecond temporal duration. Although, the use of multiple RESs for high harmonic generation (HHG) has been demonstrated experimentally with multi-cycle, linearly polarized laser pulses [30], the generation of a single dense RES capable of producing isolated attosecond pulses remains an open challenge.
@@ II Influence of drive laser waveform on TBS condition in laser-foil interaction
To enable the formation of an optimal Relativistic Electron Sheet (RES), the incident drive laser must exert a sufficiently strong ponderomotive force to expel plasma electrons from the ion background and accelerate them coherently in the forward direction. This requires the laser to generate a longitudinal push force capable of compressing electrons into a sub-skin-depth layer, ultimately overcoming the electrostatic restoring force of the ions. The necessary condition for this ”snowploughing” [12, 31] effect to enable complete electron evacuation, is that the normalized vector potential of the drive laser satisfies, $a_{0}^{d}>\omega_{p}^{2}d/c\omega_{L}$, where $\omega_{p}=\sqrt{n_{e}e^{2}/\epsilon_{0}m_{e}}$ is the plasma frequency, $d$ is the initial thickness of the plasma foil, $\omega_{L}$ is the drive laser frequency and $c$ is the speed of light [32, 33, 4, 30].
@@
Figure 1: Influence of drive pulse temporal envelope and its waveform on the formation of the electron mirror and backscattered signal. The 4040nm thick plasma slab (located from $6$ to $6.05\lambda_{L}$) is irradiated by the drive pulse possessing different waveform and temporal envelope (a) top-hat, (b) Gaussian, and (c) optimal dual-color Gaussian (with $P=0.4$) pulse, which causes various temporal evolution of electron number density. In all cases the drive beam propagates from left to right and is incident normally on the 4040nm thick pre-ionized plasma foil at $6T_{L}$. (d) compares the temporal waveforms of the single ($1\omega$) and two-color ($1\omega+2\omega$) drive Gaussian pulses with end-to-end $12T_{L}$ temporal duration and polarization along y-direction. The inset at the lower left-hand side enlarges the waveforms from $4$ to $5T_{L}$, highlighting the steepness of accelerating two-color field. The counter propagating source pulse with duration $6T_{L}$ and peak amplitude of $a_{0}=0.1$ interacts with the electron sheet and reflects partially. The squared field of the normalized TBS signals are shown in (e), indicating the generation of a single 110110as pulse (in FWHM) in case of $(1\omega+2\omega)$ drive beam. The duration of the TBS signal, resulting from interaction with the produced RES in panel (a) is 380380as. This pulse is scaled by a factor of $10$ (in blue). The reflected pulses are temporally shifted for better comparison. The corresponding frequency spectra of the TBS radiations are illustrated in panel (f). The AP generated by interaction of dual-color drive beam and foil target is composed of spectrum peaked at $26\omega_{L}$.
@@
To gain a comprehensive understanding of how the temporal waveform of the drive laser pulse influences the interaction dynamics with a dense plasma foil, we carry out PIC simulations [34]. We examine the spatio-temporal evolution of the electron number density $n_{e}(x,t)$ (normalized by critical density $n_{c}$) as shown in Figure  1(a). This figure depicts the interaction between a top-hat laser pulse—represented as a sinusoidal wave—and a nanometric (d=40nmd=40~nm) solid-density foil. The incident laser is linearly polarized (electric field along the y-axis) and propagates along the x-axis (left to right) with normalized electric field amplitude of $a_{0}^{d}=eE_{0}^{d}/mc\omega_{L}=20$ corresponding to a peak intensity of $8.6\times 10^{20}$W/cm2 for a central wave length $\lambda_{L}=800$nm. The laser pulse impinges normally on a 4040nm -thick, sharp-edged fully ionized Carbon plasma foil initially located at $x=6\lambda_{L}$, with an initial electron density of $n_{e}=21n_{c}$. This configuration ensures a relativistic near-critical density regime, characterized by the similarity parameter $S=n_{e}/a_{0}^{d}n_{c}\approx 1$ [35], which is optimal for efficient electron blowout and RES formation. Here, the critical density is defined as $n_{c}=mc\omega^{2}_{L}/e^{2}=1.7\times 10^{21}$cm-3 where $m$ denote the electron mass. The initial temperature of the plasma electrons is set as $T_{e}=100$eV. The size of each spatial grid of the simulation box is $\Delta x=\lambda_{L}/4000$, in which there are $500$ macro-particles of each species.
@@
As the sharp-fronted drive pulse reaches the target at $t=6T_{L}$, the radiation pressure at the leading edge of the pulse rapidly exceeds the electrostatic binding forces between electrons and ions. This results in a prompt and coherent expulsion of electrons in the forward direction, parallel to the laser propagation axis. The expelled electrons are compressed into a dense, ultra-thin RES that is accelerated synchronously as is evident in Figure 1(a). Although such an approach using non-adiabatic sharp rising pulse has been investigated in numerous occasions for generating RES, accessing such pulses is extremely complex and challenging. Considering the key role of pulse front steepness, plasma-based pulse front steepening was proposed in [36] and implemented in ion acceleration experiments [37]. In such experiments a separate low density target is designed prior to the main target, where relativistic non-linearities introduced by this near critical density plasma modifies the front edge of the pulse. This approach ultimately leads to a laser pulse, which is spatio-temporally compressed [38]. Once such pulses are generated, a single RES can be formed [39]. However, these methods, which are based on non-linear pulse shaping within a plasma medium, introduce a significantly higher level of experimental complexity.
@@
A typical currently available, ultrashort high power laser system delivers pulses with Gaussian temporal envelop on target. Figure 1(b) presents the electron density evolution when the same plasma foil is irradiated by such a Gaussian-shaped laser pulse with identical peak field amplitude $a_{0}^{d}=20$. Despite having the same peak intensity, $I_{0}^{d}[Wcm^{-2}]=1.37\times 10^{18}(a_{0}^{d}/\lambda_{L}[\mu m])^{2}$, the gradual rise of the Gaussian envelope fails to deliver a sufficiently strong ponderomotive force at the pulse front. As a result, the electrons are not coherently swept into a single sheet but instead experience multiple oscillations within the laser field. This leads to incomplete electron blowout, in contrast to the case presented in Figure 1(a), and the plasma surface becomes distorted and modulated, rather than forming a clean RES. The electrons remain partially bound and oscillate around the ion background, resulting in a disordered plasma structure with reduced coherence and acceleration efficiency. It is also evident in Figure 1(b), that there are signatures of partial electron ejection on the back side of the irradiated target, leading to multiple potential electron bunches of varying strength. This comparison highlights the critical role of the pulse temporal profile—particularly the sharpness of the leading edge—in determining the efficiency of RES formation. A steep front ensures a rapid onset of radiation pressure, enabling synchronized electron expulsion, while a smooth Gaussian front leads to phase mixing and multiple bunch emission leading to their structural degradation.
@@
This distinction in interaction dynamics is particularly relevant in the context of coherent Thomson backscattering for generation of as pulses. In the experiment reported by Kiefer et. al. [30], a 1010nm diamond-like carbon (DLC) was irradiated by a Gaussian-shaped drive pulse with intensity exceeding $1\times 10^{20}$Wcm-2. Due to the smooth rising edge of the Gaussian pulse, the initial few laser cycles were insufficiently intense to fully overcome the binding forces within the foil. Consequently, only a fraction of the electrons are expelled during each cycle of the laser field, resulting in the formation of multiple relativistic electron layers, or RESs, at the rear side of the target. These RESs, being only partially reflective, interact with a counter-propagating probe pulse of intensity $\sim 1\times 10^{15}$Wcm-2. As the probe pulse is back-reflected from a sequence of partially transmissive relativistic electron sheets, generating an attosecond pulse train. This process produces a reflected spectrum rich in high-order harmonics of the fundamental frequency of the probe. This mechanism stands in contrast to the interaction driven by a sharp-edged pulse front, which is capable of expelling all plasma electrons within a single half-cycle, thereby forming a single, highly reflective RES. This configuration is more favorable for generating isolated attosecond bursts.
@@
Since the RES formation is directly governed by the drive pulse field, there exists a strong correlation between the waveform of the drive pulse and the coherent electron dynamics responsible for RES generation [23, 24, 40]. Here, we demonstrate that the laser field waveform can be shaped optimally by the addition of a second harmonic component to a Gaussian drive pulse to excite plasma electron dynamics, that enables the formation and acceleration of a single, dense relativistic electron sheet. The significance of drive pulse waveform engineering has been widely explored in various high-field plasma interaction regimes, including laser wakefield acceleration [41] and attosecond XUV pulse generation via high harmonic emission from plasma surfaces [42, 43, 7, 44, 45, 46, 47]. These studies consistently highlight the importance of sub-cycle control over the laser field to manipulate electron trajectories and optimize emission characteristics.
@@
In Figure 1(c), we present the case when the drive laser waveform is shaped by mixing with the fundamental frequency $\omega_{L}$, a fraction $P$ of it’s second harmonic $2\omega_{L}$. As indicated by the arrow in Figure 1(c), this two-color linearly polarized drive configuration effectively excites plasma electrons which are compressed into a coherent structure forming a single RES. Figure 1(d) illustrates the two drive pulses employed to initiate interactions, presented in Figure 1(b-c), with different drive pulse waveforms: Gaussian (in black marked $1\omega$), and two-color Gaussian (in magenta marked $1\omega+2\omega$). However, the distinct temporal waveform of each pulse leads markedly different electron dynamics and RES formation behavior. For simulations presented in Figures 1(b) and 1(c) total energy contained within the drive laser pulse is kept constant.
@@
Figure 2: Optimal generation of single RES with waveform-controlled laser driver. (a) The longitudinal velocity of the ejected electron beam, averaged over the bunch $\beta_{x}\!>$ as a function of $P$ and $\Delta\phi$. (b) The waveform dependent maximum energy of accelerated electrons in the bunch represented by the acquired cutoff energy $\varepsilon_{c}$ in ($P,\Delta\phi$)-space. The dashed encircled region in (a) and (b) denotes the parameter space, in which a single RES is produced and accelerated. The density profile of the electron beam after traveling $5\lambda_{L}$ to the location of detector $(x=11\lambda_{L})$ are plotted for two optimal waveforms of the drive laser (c) $P=0.4$ with $\Delta\phi=0$, and (d) $P=0.6$ with $\Delta\phi=-\pi/3$, in which a single electron beam with attosecond duration is emitted. (e) shows the spatio-temporal evolution of thin foil electron density (red to yellow colormap) and drive pulse field (green to violet colormap) for a single color Gaussian driver. (f) The spacetime evolution of the laser field and electron density for single RES regime as well as the temporal evolution of phasespace parameters of electron bunches. The time axis implies the simulation time (see Fig. 1(b) and (c)) during interaction of drive beam, synthesized by (e) only fundamental frequency ($1\omega$) or (f) by combination of in phase frequencies $(1\omega+2\omega)$ with energy ratio $P=0.4$. The dashed lines in (e) and (f) represent trajectory of the representative electrons which are subject to the normalized drive field $B_{z}$. (g) and (h) show temporal variation of the longitudinal and transverse components of the velocity of the representative electrons (black curve for $1\omega$ and magenta for $1\omega+2\omega$ field). The temporal evolution of (i) bunch peak density and (j) thickness during the interaction with a counter propagating source pulse with the RES generated by the two-color Gaussian driver beam (with $P=0.4$ and $\Delta\phi=0$). The FWHM of the bunch is calculated by fitting a Gaussian function on the electron density profile. Greater longitudinal velocity of RES during interaction phase accounts for larger compression of the AP in case of synthesized pulse.
Once a single forward propagating RES is formed, a cross-polarized, counter-propagating source pulse is introduced at a carefully timed delay relative to the drive pulse. This source pulse interacts with the RES at normal incidence and is reflected, undergoing a Doppler frequency upshift and temporal compression into a single attosecond pulse (AP). The efficacy of RES for providing optimal conditions for TBS of a linearly polarized Gaussian source pulse (electric field along the z-axis), with normalized amplitude with $a_{0}=0.1$ and a temporal duration of 44fs (corresponding to $1.5T_{L}$ in FWHM of intensity) is summarized in Figure 1(e-f). This correlation between the quality of RES and back reflected pulse, is quantitatively illustrated in Figure 1(e), which compares the temporal profiles of TBS pulses generated by RESs formed under different drive pulse temporal waveforms. The two-color drive pulse produces a compressed AP with a duration of (intensity FWHM) $110$ as and a peak intensity of $\approx 2.6\times 10^{15}~Wcm^{-2}$ (corresponding to $a_{0}^{2}\approx 1.2\times 10^{-3}$), significantly outperforming the weaker and longer AP (380380as) generated by even the ideal top-hat pulse. This clearly demonstrates the enhanced efficiency and coherence enabled by waveform shaping. The temporal profile for the Gaussian case does not lead to single AP formation.
The corresponding frequency spectra presented in Figure 1(f) further reinforce this conclusion. The ideally top-hat drive pulse is plotted in blue curve and bicolor Gaussian pulse in magenta curve. The RES formed by the two-color drive pulse reaches a higher velocity, resulting in a larger central frequency shift of the reflected spectrum up to $26\omega_{L}$ and a broader bandwidth compared to the sharp rising pulse front case, ideal for shorter AP. In contrast, in case of Gaussian drive (black curve in Figure 1(f)), most of the scattered energy is concentrated in low order harmonics induced in the reflected pulse, in line with the observation discussed in Figure 1(b).
These results underscore the critical role of multi-frequency waveform synthesis in controlling RES formation and optimizing attosecond pulse generation. A deeper analysis of this mechanism is presented in Section III, where the dynamics of single RES formation and its impact on coherent backscattering are discussed in detail. Once single RES is formed, it can be combined with double-layer target configuration [48, 49, 50], to boost the longitudinal Lorentz factor $\gamma_{x}=(1-\beta_{x}^{2})^{-1/2}$ further and enhance the backreflected as pulse, where $\beta_{x}$ is the longitudinal component of the RES velocity in units of $c$.